OpenStudy (dtan5457):

Solve for sin a/2 when cos a=4/5 , o

OpenStudy (dtan5457):

just to check i got \[\frac{ 3\sqrt{10} }{ 10 }\]

OpenStudy (dtan5457):

correct?

OpenStudy (dtan5457):

and wouldn't cos/2 be the same thing?

OpenStudy (welshfella):

I guess you would use the identity cos a = 1 - 2 sin^2 (a/2) ?

OpenStudy (welshfella):

how did you get the 3?

OpenStudy (dtan5457):

I just used the half angle formula

OpenStudy (welshfella):

oh - i cant remember that one lol

OpenStudy (welshfella):

I got sqrt10/10 using my formula. Again its a while since I did trig....

OpenStudy (dtan5457):

@Nnesha

OpenStudy (welshfella):

cosa = 1 - 2 sin^2 (a/2) 4/5 = 1 - 2 * 10/100 4/5 = 1 - 1/5

OpenStudy (welshfella):

so it does fit

OpenStudy (dtan5457):

ohhh, your saying that the answer is sqrt of 1/10?

OpenStudy (dtan5457):

if so , I see what i did wrong, i'm suppose to subtract not add 5+4

OpenStudy (dtan5457):

so cos/2 would be \[\frac{ 3\sqrt{10} }{ 10 }\] sin/2=\[\sqrt{\frac{ 1 }{ 10 }}\]

OpenStudy (dtan5457):

@welshfella

OpenStudy (welshfella):

yes

OpenStudy (dtan5457):

so both those are correct?

OpenStudy (welshfella):

yes

OpenStudy (dtan5457):

now i need to solve for tan /2

OpenStudy (dtan5457):

im actually getting just 1/3

OpenStudy (welshfella):

oh - now i definitely cant remember the tan/2 formula , i'm afraid

OpenStudy (dtan5457):

looks a lot easier actually

OpenStudy (dtan5457):

im just getting (3/5)/1+4/5=1/3

OpenStudy (welshfella):

yea - they are vaguely familiar lol

OpenStudy (dtan5457):

are you able to do it?

OpenStudy (dtan5457):

@TheSmartOne

OpenStudy (welshfella):

yes 1/3 is correct

OpenStudy (dtan5457):

ok ty